# STAR-GO: Improving Protein Function Prediction by Learning to Hierarchically Integrate Ontology-Informed Semantic Embeddings

### v1

@@ Abstract
Availability: Code and pre-trained models are available at https://github.com/boun-tabi-lifelu/stargo.
@@ 1 Introduction
In line with advances in GO embedding techniques, protein function prediction models that incorporate GO embeddings into their architectures have also been recently proposed. DeepGOZero [12] and its successor DeepGO-SE [14] employ EL Embeddings, representing GO terms as geometric objects (n-dimensional balls) constrained by logical axioms. This formulation encodes the ontology’s formal relationships, enabling zero-shot prediction for unseen GO terms. DeepGO-SE further introduces approximate semantic entailment by ensembling models built on pretrained protein language models [20], yet its representations remain fundamentally axiom-based. PFresGO [17] instead employs anc2vec [5] embeddings and integrates them with protein representations via cross-attention. While it captures hierarchical and membership relations, anc2Vec encodes GO terms as one-hot vectors, which limits its zero-shot capability. Although these studies benefit from incorporating ontological structure into protein function prediction, their GO embedding techniques either prioritize graph structure at the expense of the semantics of term descriptions or capture semantic content but fail to integrate the ontology’s hierarchy.
@@ 2.1 Dataset
We train and evaluate STAR-GO on a dataset curated from DeepFRI [10], which is commonly used in the protein function prediction literature. The dataset comprises $36,641$ protein sequences with coverage of $2,752$ GO terms, including the following subontology terms: $489$ Molecular Function, $1,943$ Biological Process, and $320$ Cellular Component, where each GO term is linked to at least $50$ non-redundant Protein Data Bank (PDB) chains. The dataset is split into training (80%; 29,902 sequences), validation (10%; 3,323 sequences), and test (10%; 3,416 sequences) sets (see Supplementary Materials Table 1 for details). The test set consists only of proteins with at least one experimentally validated functional annotation in each of the three GO categories, with protein chain lengths limited to 1,000 residues. To ensure non-redundancy between the training and test sets, cd-hit [16] was applied using 95% sequence identity threshold.
For zero-shot evaluation, we follow DeepGOZero’s protocol [12], using the same similarity-based split of the UniProt/Swiss-Prot Knowledgebase (version 2021_04, on September 29, 2021; see Supplementary Table 2).We hold out the same 16 GO terms as DeepGOZero (5 MF, 7 BP, 4 CC), which are selected from classes with >100>100 annotations. We remove all protein-class associations for these terms from the training set before applying the true path rule, which propagates annotations of GO terms to their ancestors, ensuring hierarchical consistency. We use the basic version of GO with the true path rule applied for training and evaluations, specifically the June 2020 release [^1] https://release.geneontology.org/2020-06-01/ontology/go-basic.obo in the standard setting and the November 2021 release [^2] https://release.geneontology.org/2021-11-16/ontology/go-basic.obo in the zero-shot setting.
@@ 2.2 STAR-GO
The STAR-GO model adopts an autoencoder architecture, given in Figure 1, where protein and GO embeddings are fused within a unified transformer framework. The encoder takes protein embeddings obtained from a pretrained language model, while the decoder predicts GO term probabilities based on the fusion of the protein and GO representations via cross-attention. To ensure consistency with the hierarchical structure of the GO ontology, the decoder processes the GO embeddings arranged in a topological order from ancestors to descendants. Consequently, the proposed hierarchical decoding framework captures the dependencies across multiple functional levels.
@@ 2.4 Transformer module
The proposed module learns to associate protein sequences with GO terms by jointly processing residue-level embeddings representing protein sequences and term-level embeddings representing GO semantics and hierarchy. The architecture follows an autoencoder design where cross-attention combines protein and GO representations. The encoder consists of $N_{\text{enc}}$ self-attention layers that contextualize residue embeddings, capturing dependencies along the sequence. The decoder comprises $N_{\text{dec}}$ layers that combine self-attention and cross-attention operations. The decoder’s self-attention operates over GO terms with a causal mask that respects their topological order, allowing information to flow from general to specific functions per the GO hierarchy. Cross-attention layers link the GO term representations with the residue embeddings, allowing each term to attend to the most informative residues relevant to its semantics. The model is trained for the supervised function prediction task, encouraging the model to capture task-relevant associations between residue-level and GO-level representations.
2.4.1 Encoder
@@
Self-Attention Layers. The encoder refines the projected residue embeddings through $N_{\text{enc}}$ stacked self-attention layers. Let $\mathbf{H}^{(l)}_{\text{enc}}\in\mathbb{R}^{L\times d}$ denote the hidden representation of the protein sequence at the $l$-th encoder layer, where $\mathbf{H}^{(0)}_{\text{enc}}$ is the input defined above. Each layer updates these representations as follows:
where $\mathrm{MHA}$ denotes multi-head attention [23], $\mathrm{FFN}$ is a position-wise feed-forward network, and $\mathrm{LN}$ denotes layer normalization. $\tilde{\mathbf{M}}_{\text{seq}}$ is an additive mask for sequence padding. After $N_{\text{enc}}$ layers, the encoder outputs refined residue embeddings $\mathbf{H}^{(N_{\text{enc}})}_{\text{enc}}$.
@@ 2.4.2 Decoder
Input Representation and Ordering. Each GO term within a subontology is represented by the embeddings obtained from the GO embedding module, $\mathbf{E}\in\mathbb{R}^{T\times d_{\text{go}}}$, where $T$ denotes the number of GO terms. To preserve the hierarchical dependencies among GO terms, they are arranged in a topological order $(\mathrm{GO}_{\pi(1)},\ldots,\mathrm{GO}_{\pi(T)})$, where $\pi$ defines a breadth-first traversal from root to leaves. This ordering enables information to propagate along parent–child relationships during decoding. The GO embeddings are then projected into the decoder’s latent space, producing the initial representation $\mathbf{H}^{(0)}_{\text{dec}}$:
@@
Self-Attention and Cross-Attention. The decoder consists of $N_{\text{dec}}$ layers. At each layer, causal self-attention is first applied over GO terms using a lower-triangular mask $\mathbf{C}$, ensuring that each term attends only to its ancestors. Subsequently, cross-attention incorporates the encoded protein representations, allowing GO terms to attend to residue-level features relevant to their semantics. Thus, each GO term representation aggregates information from both its ancestors and the relevant residues of the protein. A feed-forward transformation with residual connections and layer normalization refines the decoder’s final representations. Formally, let $\mathbf{H}^{(l)}_{\text{dec}}\in\mathbb{R}^{T\times d}$ denote the hidden representation of the GO terms at layer $l$, then each layer is computed as:
$\textbf{Self-attention:}\quad\hat{\mathbf{H}}^{(l)}_{\text{dec}}$ $=\mathrm{MHA}\!\left(\mathbf{H}^{(l-1)}_{\text{dec}},\mathbf{H}^{(l-1)}_{\text{dec}};\tilde{\mathbf{M}}_{\text{dec}}\right),$ (6)
$\textbf{Cross-attention:}\quad\tilde{\mathbf{H}}^{(l)}_{\text{dec}}$ $=\mathrm{MHA}\!\left(\mathbf{S}^{(l)}_{\text{dec}},\mathbf{H}^{(N_{\text{enc}})}_{\text{enc}};\mathbf{0}\right),$ (8)
$\textbf{Feed-forward:}\quad\mathbf{H}^{(l)}_{\text{dec}}$ $=\mathrm{LN}\!\left(\mathbf{C}^{(l)}_{\text{dec}}+\mathrm{FFN}\!\left(\mathbf{C}^{(l)}_{\text{dec}}\right)\right)$. (10)
@@ 2.4.3 Prediction Head and Objective
Output Mapping. For each GO term $t$, the final decoder representation is mapped to a logit $z_{t}$ through a two-layer feed-forward projection:
where $\hat{y}_{t}$ denotes the predicted probability that the protein is associated with GO term $t$, and $\sigma(\cdot)$ is the sigmoid activation function that maps each logit $z_{t}$ to a probability value in the range $(0,1)$.
Training Objective. Given binary ground-truth $\mathbf{y}\in\{0,1\}^{T}$, where $y_{t}=1$ if the protein is annotated with GO term $t$ and 0 otherwise, the model is trained using binary cross-entropy:
During training, only the projection layers, Transformer blocks, and prediction head are optimized, while the pretrained protein and GO embedding modules remain frozen.
@@ 3 Results
We compared STAR-GO with state-of-the-art models to evaluate its protein function prediction performance and generalization to unseen GO terms. We further examined the contribution of the STAR-GO components with ablation studies. Performance was evaluated using $F_{max}$ for protein-centric accuracy and Macro-AUPR and AUC for term-centric generalization across GO terms. Implementation details, baselines, and evaluation metrics are provided in the Supplementary Materials.
@@ 3.1 STAR-GO achieves competitive performance
We compared STAR-GO with state-of-the-art baselines, including PFresGO, DeepFRI, TALE+, DeepGOZero, and DeepGO-SE, using $F_{\text{max}}$, Macro AUPR, and AUC metrics (Table 1). Across all ontologies, STAR-GO exhibits competitive performance, matching or exceeding most baselines. In BP, DeepGO-SE attains the highest overall $F_{\text{max}}$ and Macro AUPR, while STAR-GO achieves the best AUC (0.989). Within CC ontology, STAR-GO surpasses all baselines in Macro AUPR and AUC while maintaining comparable $F_{\text{max}}$ to PFresGO. For MF, STAR-GO performs on par with DeepGO-SE, achieving close $F_{\text{max}}$ and AUC values. STAR-GO consistently ranks among the top-performing methods and achieves the strongest AUC values across all ontologies, indicating strong term-level discriminability.
@@ 3.2 STAR-GO generalizes to unseen GO terms
For zero-shot evaluation, we compared STAR-GO with DeepGOZero and DeepGO-SE, which have zero-shot capabilities, using the zero-shot dataset, where selected GO terms were excluded from all protein–term associations. To assess the contributions of different GO information sources, we ablated the structural and textual components of our GO embedding module and retrained the resulting variants, denoted STARS (structural only, using anc2vec embeddings) and START (textual-only, using SBERT-BioBERT embeddings), and STARST model combining both representations. All models, including baselines, were trained on the zero-shot dataset and evaluated on the full test set, with term-specific AUCs reported for each held-out GO term in Table 2.
Across the 16 held-out GO terms, STAR-GO variants collectively achieve the highest zero-shot AUCs in 13 cases, consistently outperforming DeepGOZero and DeepGO-SE. Performance varies by ontology and GO embeddings: START achieves the best results on most Molecular Function and Biological Process terms (e.g., GO:0001228 = 0.949, GO:0046330 = 0.960), highlighting the strength of semantic information captured from textual definitions. STARS performs best on a few terms (e.g., GO:0032760 = 0.930, GO:1904813 = 0.903) but performs poorly for MF terms. The combined STARST model maintains stable and competitive performance across all ontologies, closely matching the textual variant.
An independent subsumption-prediction evaluation of the GO embeddings further contextualized these results (see Supplementary Materials). The anc2vec performed the best in recovering hierarchical parent–child relations, confirming its strong topological representation of GO. However, anc2vec embeddings generalized less effectively to unseen terms, particularly for MF terms, resulting in the lower zero-shot AUCs of STARS. Conversely, SBERT-BioBERT embeddings performed worse in subsumption prediction but generalized better to unseen terms, indicating that semantic representations transfer more effectively to novel ontology concepts. The integrated STARST model reflects this complementarity, achieving balanced performance across ontologies by leveraging both semantic and structural information. While STARS requires retraining when new GO terms are added due to the fixed anc2vec embeddings, START and STARST can generate embeddings for novel terms directly from textual definitions, enabling continuous adaptation to ontology updates without retraining.
@@
Table 2: Zero-shot performance of STAR-GO with unseen GO term AUCs. The zero-shot columns denote that models were optimized on an ablated dataset excluding the reported GO terms from all protein–term associations. Trained columns reports the results for the corresponding protein-term pairs included during training. STAR variants use different GO embeddings: $S$ (structural, anc2Vec), $T$ (textual, SBERT-BioBERT), and STST (combined, proposed in this work). The best AUC within each section is shown in bold.
Zero-shot Trained
@@
Ontology Term STARST START STARS DeepGOZero DeepGO-SE STARST DeepGOZero DeepGO-SE
MF GO:0001227 0.891 0.944 0.557 0.257 0.759 0.955 0.932 0.952
GO:0001228 0.938 0.949 0.508 0.574 0.791 0.961 0.948 0.961
GO:0003735 0.923 0.915 0.301 0.400 0.066 0.994 0.940 0.995
GO:0004867 0.742 0.884 0.807 0.972 0.568 0.955 0.985 0.992
GO:0005096 0.904 0.907 0.750 0.847 0.517 0.962 0.938 0.735
BP GO:0000381 0.973 0.988 0.957 0.855 0.751 0.987 0.906 0.992
GO:0032729 0.895 0.895 0.948 0.870 0.982 0.893 0.932 0.996
GO:0032755 0.895 0.889 0.957 0.719 0.966 0.939 0.884 0.979
GO:0032760 0.900 0.883 0.930 0.861 0.803 0.950 0.925 0.821
GO:0046330 0.949 0.960 0.923 0.855 0.866 0.966 0.904 0.888
GO:0051897 0.893 0.942 0.913 0.772 0.929 0.962 0.888 0.957
GO:0120162 0.734 0.779 0.811 0.637 0.491 0.817 0.738 0.957
$C$ GO:0005762 0.998 0.998 0.991 0.889 0.600 0.999 0.874 0.995
GO:0022625 0.989 0.937 0.950 0.898 0.487 0.995 0.893 0.991
GO:0042788 0.940 0.963 0.930 0.858 0.532 0.990 0.889 0.990
GO:1904813 0.675 0.806 0.903 0.653 0.658 0.895 0.792 0.955

### v2

@@ Abstract
Availability: Code and pre-trained models are available at https://github.com/boun-tabi-lifelu/stargo
(https://doi.org/10.5281/zenodo.18643082)
footnotetext: This article has been accepted for publication in Bioinformatics, Published by Oxford University Press.
@@ 1 Introduction
In line with advances in GO embedding techniques, protein function prediction models that incorporate GO embeddings into their architectures have also been recently proposed. DeepGOZero [14] and its successor DeepGO-SE [16] employ EL Embeddings, representing GO terms as geometric objects (n-dimensional balls) constrained by logical axioms. This formulation encodes the ontology’s formal relationships, enabling zero-shot prediction for unseen GO terms. DeepGO-SE further introduces approximate semantic entailment by ensembling models built on pretrained protein language models [23], yet its representations remain fundamentally axiom-based. PFresGO [20] instead employs anc2vec [7] embeddings and integrates them with protein representations via cross-attention. While it captures hierarchical and membership relations, anc2Vec encodes GO terms as one-hot vectors, which limits its zero-shot capability. TransFew [4] combines BioBERT-derived textual embeddings with GCN-encoded hierarchical relationships and fuses them with protein representations via cross-attention, targeting rare GO term prediction. Although these studies benefit from incorporating ontological structure into protein function prediction, none jointly optimizes semantic and structural GO embeddings within a hierarchical decoding framework that captures functional dependencies across ontology levels.
@@ 2.1 Dataset
We train and evaluate STAR-GO on a dataset curated from DeepFRI [12], which is commonly used in the protein function prediction literature. The dataset comprises $36,641$ protein sequences with coverage of $2,752$ GO terms, including the following subontology terms: $489$ Molecular Function, $1,943$ Biological Process, and $320$ Cellular Component, where each GO term is linked to at least $50$ non-redundant Protein Data Bank (PDB) chains.
The dataset is split into training (80%; 29,902 sequences), validation (10%; 3,323 sequences), and test (10%; 3,416 sequences) sets (see Supplementary Table 1 for details). The test set consists only of proteins with at least one experimentally validated functional annotation in each of the three GO categories, with protein chain lengths limited to 1,000 residues. To ensure non-redundancy between the training and test sets, cd-hit [18] was applied using 95% sequence identity threshold.
For zero-shot evaluation, we follow DeepGOZero’s protocol [14], using the same similarity-based split of the UniProt/Swiss-Prot Knowledgebase (version 2021_04, on September 29, 2021; see Supplementary Table 2).We hold out the same 16 GO terms as DeepGOZero (5 MF, 7 BP, 4 CC), which are selected from classes with >100>100 annotations. We remove all protein-class associations for these terms from the training set before applying the true path rule, which propagates annotations of GO terms to their ancestors, ensuring hierarchical consistency.
We use the basic version of GO with the true path rule applied for training and evaluations, specifically the June 2020 release [^1] https://release.geneontology.org/2020-06-01/ontology/go-basic.obo in the standard setting and the November 2021 release [^2] https://release.geneontology.org/2021-11-16/ontology/go-basic.obo in the zero-shot setting.
@@ 2.2 STAR-GO
STAR-GO adopts an encoder-decoder architecture, given in Figure 1, where protein and GO embeddings are fused within a unified Transformer framework. The encoder takes protein embeddings obtained from a pretrained language model, while the decoder predicts GO term probabilities based on the fusion of the protein and GO representations via cross-attention. To ensure consistency with the hierarchical structure of the GO ontology, the decoder processes the GO embeddings arranged in a topological order from ancestors to descendants. Consequently, the proposed hierarchical decoding framework captures the dependencies across multiple functional levels.
@@ 2.4 Transformer module
The proposed module learns to associate protein sequences with GO terms by jointly processing residue-level embeddings representing protein sequences and term-level embeddings representing GO semantics and hierarchy. The architecture follows an encoder-decoder design where cross-attention combines protein and GO representations. The encoder consists of $N_{\text{enc}}$ self-attention layers that contextualize residue embeddings, capturing dependencies along the sequence. The decoder comprises $N_{\text{dec}}$ layers that combine self-attention and cross-attention operations. The decoder’s self-attention operates over GO terms with a causal mask that respects their topological order, allowing information to flow from general to specific functions per the GO hierarchy. Cross-attention layers link the GO term representations with the residue embeddings, allowing each term to attend to the most informative residues relevant to its semantics. We adopt cross-attention as the fusion mechanism because it has been successfully applied in prior protein function prediction models, including PFresGO and TransFew, to integrate protein and GO term representations. The model is trained for the supervised function prediction task, encouraging the model to capture task-relevant associations between residue-level and GO-level representations. Our implementation builds on HuggingFace Transformers [27] encoder and decoder layers. For a formal definition of the architecture, please see Supplementary Materials.
@@
Self-Attention Layers. The encoder refines the projected residue embeddings through $N_{\text{enc}}$ stacked self-attention layers, each comprising multi-head attention [26], a position-wise feed-forward network, and layer normalization with residual connections. After $N_{\text{enc}}$ layers, the encoder outputs refined residue embeddings $\mathbf{H}^{(N_{\text{enc}})}_{\text{enc}}$.
@@ 2.4.1 Decoder
Input Representation and Ordering. Each GO term within a subontology is represented by the embeddings obtained from the GO embedding module, $\mathbf{E}\in\mathbb{R}^{T\times d_{\text{go}}}$, where $T$ denotes the number of GO terms. To preserve the hierarchical dependencies among GO terms, they are arranged in a topological order $(\mathrm{GO}_{\pi(1)},\ldots,\mathrm{GO}_{\pi(T)})$, where $\pi$ defines a breadth-first traversal from root to leaves. This ordering enables information to propagate along parent-child relationships during decoding. The GO embeddings are then projected into the decoder’s latent space, producing the initial representation $\mathbf{H}^{(0)}_{\text{dec}}$:
@@
Self-Attention and Cross-Attention. The decoder consists of $N_{\text{dec}}$ layers, each containing causal self-attention, cross-attention, and a feed-forward sub-layer with residual connections and layer normalization. Causal self-attention is applied over GO terms using a lower-triangular mask $\tilde{\mathbf{M}}_{\text{dec}}$, ensuring that each term attends only to terms preceding it in topological order, providing an inductive bias aligning with GO hierarchy. Subsequently, cross-attention incorporates the encoded protein representations $\mathbf{H}^{(N_{\text{enc}})}_{\text{enc}}$, allowing GO terms to attend to residue-level features relevant to their semantics. Thus, each GO term representation aggregates information from both its ancestors and the relevant residues of the protein.
@@ 2.4.2 Prediction Head and Objective
The final decoder output is mapped through a two-layer feed-forward projection with GELU activation and sigmoid to get the predicted probability that the protein is associated with each GO term $t$. The model is trained using binary cross-entropy. During training, only the projection layers, Transformer blocks, and prediction head are optimized, while the pretrained protein and GO embedding modules remain frozen.
@@ 3 Results
We compared STAR-GO with state-of-the-art models to evaluate its protein function prediction performance and generalization to unseen GO terms. We further examined the contribution of STAR-GO components with ablation studies. Performance was evaluated using $F_{max}$ for protein-centric accuracy and Macro AUPR and AUC for term-centric generalization across GO terms. Implementation details, baselines, and evaluation metrics are provided in the Supplementary Materials.
@@ 3.1 STAR-GO achieves competitive performance
We compared STAR-GO with state-of-the-art baselines, including PFresGO, DeepFRI, TALE+, DeepGOZero, DeepGO-SE, and TransFew, using $F_{\text{max}}$, Macro AUPR, and AUC metrics (Table 1). Across all ontologies, STAR-GO exhibits competitive performance, matching or exceeding most baselines. In BP, DeepGO-SE attains the highest overall $F_{\text{max}}$ and Macro AUPR, while STAR-GO achieves the best AUC (0.989). Within CC ontology, STAR-GO surpasses all baselines in Macro AUPR and AUC while maintaining comparable $F_{\text{max}}$ to PFresGO. For MF, STAR-GO performs on par with DeepGO-SE, achieving close $F_{\text{max}}$ and AUC values. STAR-GO consistently ranks among the top-performing methods and achieves the strongest AUC values across all ontologies, indicating strong term-level discriminability.
@@ 3.2 STAR-GO generalizes to unseen GO terms
Table 2: Zero-shot performance of STAR-GO with unseen GO term AUCs. The zero-shot columns denote that models were optimized on an ablated dataset excluding the reported GO terms from all protein–term associations. Supervised columns reports the results for the corresponding protein-term pairs included during training. STAR variants use different GO embeddings: $S$ (structural, anc2Vec), $T$ (textual, SBERT-BioBERT), and STST (combined, proposed in this work). The best AUC within each section is shown in bold.
Zero-shot Supervised
@@
Ontology Term STARST START STARS DeepGOZero DeepGO-SE TransFew STARST DeepGOZero DeepGO-SE
MF GO:0001227 0.891 0.944 0.557 0.257 0.759 0.395 0.955 0.932 0.952
GO:0001228 0.938 0.949 0.508 0.574 0.791 0.533 0.961 0.948 0.961
GO:0003735 0.923 0.915 0.301 0.400 0.066 0.459 0.994 0.940 0.995
GO:0004867 0.742 0.884 0.807 0.972 0.568 0.175 0.955 0.985 0.992
GO:0005096 0.904 0.907 0.750 0.847 0.517 0.359 0.962 0.938 0.735
BP GO:0000381 0.973 0.988 0.957 0.855 0.751 0.423 0.987 0.906 0.992
GO:0032729 0.895 0.895 0.948 0.870 0.982 0.866 0.893 0.932 0.996
GO:0032755 0.895 0.889 0.957 0.719 0.966 0.754 0.939 0.884 0.979
GO:0032760 0.900 0.883 0.930 0.861 0.803 0.691 0.950 0.925 0.821
GO:0046330 0.949 0.960 0.923 0.855 0.866 0.496 0.966 0.904 0.888
GO:0051897 0.893 0.942 0.913 0.772 0.929 0.683 0.962 0.888 0.957
GO:0120162 0.734 0.779 0.811 0.637 0.491 0.600 0.817 0.738 0.957
$C$ GO:0005762 0.998 0.998 0.991 0.889 0.600 0.090 0.999 0.874 0.995
GO:0022625 0.989 0.937 0.950 0.898 0.487 0.397 0.995 0.893 0.991
GO:0042788 0.940 0.963 0.930 0.858 0.532 0.520 0.990 0.889 0.990
GO:1904813 0.675 0.806 0.903 0.653 0.658 0.724 0.895 0.792 0.955
For zero-shot evaluation, we compared STAR-GO with DeepGOZero, DeepGO-SE, and TransFew, which have zero-shot capabilities, using the zero-shot dataset, where selected GO terms were excluded from all protein–term associations. To assess the contributions of different GO information sources, we ablated the structural and textual components of our GO embedding module, denoted STARS (structural only, using anc2vec embeddings) and START (textual-only, using SBERT-BioBERT embeddings), and STARST model combining both representations. All models, including baselines, were trained on the zero-shot dataset and evaluated on the full test set, with term-specific AUCs reported for held-out GO terms in Table 2.
Across the 16 held-out GO terms, STAR-GO variants collectively achieve the highest zero-shot AUCs in 13 cases, consistently outperforming DeepGOZero, DeepGO-SE and TransFew. Performance varies by ontology and GO embeddings: START achieves the best results on most Molecular Function and Biological Process terms (e.g., GO:0001228 = 0.949, GO:0046330 = 0.960), highlighting the strength of semantic information captured from textual definitions. STARS performs best on a few terms (e.g., GO:0032760 = 0.930, GO:1904813 = 0.903) but performs poorly for MF terms. The combined STARST model maintains stable and competitive performance across all ontologies, closely matching the textual variant.
An independent subsumption-prediction evaluation of the GO embeddings further contextualized these results (see Supplementary Table S3). The anc2vec performed the best in recovering hierarchical parent–child relations, confirming its strong topological representation of GO. However, anc2vec embeddings generalized less effectively to unseen terms, particularly for MF terms, resulting in the lower zero-shot AUCs of STARS. Conversely, SBERT-BioBERT embeddings performed worse in subsumption prediction but generalized better to unseen terms, indicating that semantic representations transfer more effectively to novel ontology concepts. The integrated STARST model reflects this complementarity, achieving balanced performance across ontologies by leveraging both semantic and structural information. While STARS requires retraining when new GO terms are added due to the fixed anc2vec embeddings, START and STARST can generate embeddings for novel terms directly from textual definitions, enabling continuous adaptation to ontology updates without retraining. A seed sensitivity analysis over five random seeds confirmed the stability of these zero-shot results, with low standard deviations across all held-out terms (see Supplementary Table S5 and Table S6).